OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

just go to wolfram alpha, and plug in the equation. it will give u answers with steps and graphs if u need them!

OpenStudy (anonymous):

it is my lifesaver!

OpenStudy (anonymous):

4(x-6)=-2(x-3) 4x-24=2x-6 +6 +6 4x-18=2x /2 /2 2x-9=x +9 +9 2x=x+9 -x -x 3x=9 /3 /3 x = 3